# Semantic-Aware Resource Management for C-V2X Platooning via Multi-Agent Reinforcement Learning

### v1

@@ II-B Novel Metrics for Resource Management in platooning Networks
In vehicular platooning networks, effective resource management is crucial for maintaining efficient communication, particularly for autonomous driving applications. To optimize resource allocation and improve overall system performance, we propose novel metrics based on semantic entropy and quality of experience (QoE). These metrics aim to assess the efficiency of semantic communication by directly quantifying the amount of meaningful information being transmitted, going beyond traditional metrics that focus mainly on bit-level accuracy.
Unlike Shannon entropy, which measures the uncertainty of a source in terms of bits, semantic entropy quantifies the amount of meaningful information that a source carries with respect to a specific task. Calculating exact semantic entropy can be intractable for complex tasks. As such, we adopt deep learning (DL) models, such as DeepSC for single-modal and MU-DeepSC for multi-modal tasks, to approximate the semantic entropy as described in [67]. This approximation, expressed in semantic units (suts), serves as the basis for optimizing resource allocation in platooning networks, where the goal is to transmit only the most critical information.
We introduce a QoE model that assesses the overall performance of semantic communication systems by combining semantic accuracy and semantic rate. The semantic accuracy measures the correctness of the transmitted information, while the semantic rate indicates the amount of semantic information delivered per second. Both metrics are essential in vehicular networks, where communication must be both accurate and timely to ensure safe and coordinated actions within a platoon.
For single-modal text communication, We employ semantic similarity [65] as a performance metric (i.e., denoted as ${\xi_{SM}}$) to evaluate the accuracy of semantic transmission. This metric quantifies how closely the received message resembles the transmitted message in terms of its meaning. The semantic similarity can be expressed as:
@@
where $0\leq{\xi_{SM}}\leq 1$, with ${\xi_{SM}}=1$ indicating the highest similarity between the two sentences and ${\xi_{SM}}=0$ implying no similarity. The term ${u_{m}^{T}}$ represents the average number of transmitted semantic symbols by vehicle $m$, and $SIN{R_{T,n,m/BS}}\left[k\right]$ denotes the SINR for text communication received at vehicle $m$ in platoon $p$ or at the BS over subchannel $k$. The function ${\Psi_{SM}}$ maps these inputs to the semantic similarity score. For multi-modal communication, such as the fusion of text and image data, the joint performance of both transmissions is assessed using semantic similarity, referred to as answer accuracy $\xi_{MM}$. This metric evaluates how well the combined text and image information matches the expected outcome. The semantic similarity in this context is given by:
@@
where ${u_{m}^{T}}$ and ${u_{m}^{I}}$ denote the average number of transmitted semantic symbols for text and image, respectively, by vehicle $m$. The function ${\Psi_{MM}}$ maps these inputs to the overall semantic similarity or answer accuracy score.
The semantic rate measures the volume of meaningful information transmitted by a vehicle per second, reflecting the effectiveness of data communication in vehicular networks. Unlike traditional data rates, that focus on raw data transfer, the semantic rate emphasizes the transmission of information that is relevant for understanding and decision-making processes in specific tasks. The semantic rate ${\phi_{n}}$ is defined as:
${\phi_{n}}=\frac{{W{{\tilde{H}}_{SM}}}}{{u_{m}^{T}}},$ (15)
where $W$ represents the bandwidth allocated to the communication channel, ${{{\tilde{H}}_{SM}}}$ represents the approximate semantic entropy of the textual data. Semantic entropy is a critical concept that measures the amount of meaningful content within a message that is necessary for a specific task, rather than merely quantifying the raw data. Drawing from the work of [68], semantic entropy $H(X;Y)$ is defined as the minimum expected number of semantic symbols required to accurately predict or infer a task $Y$ from the data $X$.
$\begin{array}[]{l}H(X;Y)=\mathop{\min}\limits_{{{\rm E}_{s}}}{\rm E}(\dim(Cod{e^{{{\rm E}_{s}}}}(X)))\\ {\rm{}}\qquad\qquad\approx\mathop{\min}\limits_{s{e_{\alpha}}}{\rm E}(\dim(Cod{e^{s{e_{\alpha}}}}(X)))\\ s.t.\qquad{\rm{}}P(Y|Cod{e^{s{e_{\alpha}}}}(X)){\rm{}}\approx P(Y|X),\end{array}$ (16)
where $Cod{e^{{{\rm E}_{s}}}}(X)$ represents the semantic symbols extracted from $X$ using a semantic encoder ${{{\rm E}_{s}}}$, and $P(Y|Cod{e^{s{e_{\alpha}}}}(X))$ approximates the conditional probability of predicting $Y$ given $X$. This definition emphasizes that the goal of semantic entropy is to retain only the most relevant information for a particular task, optimizing communication efficiency, especially in dynamic and resource-constrained environments like vehicular networks. Computing an exact measure of semantic entropy is challenging due to the difficulty in determining an optimal semantic encoder. To address this, we use a deep learning-based approximation denoted as ${{{\tilde{H}}_{SM}}}$. This approach allows us to estimate the semantic content of a message by employing a well-trained semantic encoder ${s{e_{\alpha}}}$, capturing the key elements of the original data $X$ that are relevant for predicting the task $Y$.
@@
${\phi_{{n_{{}_{T}}}}}=\frac{{W\tilde{H}_{MM}^{T}}}{{u_{m}^{T}}},{\rm{}}{\phi_{{n_{{}_{I}}}}}=\frac{{W\tilde{H}_{MM}^{I}}}{{u_{m}^{T}}},$ (17)
@@
$\sum_{t=1}^{T}\sum_{m=1}^{M}{{\beta_{n,k}}}\phi_{k}^{d}[m,t]\geq\frac{B_{s}}{\Delta T},\quad k\in K,$ (18)
where $B_{s}$ denotes the size of the periodically generated V2V payload in semantic units (suts), $\Delta T$ is the channel coherence time, and $\phi_{k}^{d}[m,t]$ represents the semantic rate of the $k$-th V2V link for vehicle $m$ at time slot $t$. The variable ${{\beta_{n,k}}}$ indicates whether vehicle $m$ is assigned to subchannel $k$ (i.e., ${{\beta_{n,k}}}=1$ if assigned, 0 otherwise). This condition ensures that the required amount of semantic information is delivered within the time constraints, thus enabling safe and synchronized movements within the platoon.
Different vehicles within a platoon may prioritize these different metrics depending on their specific application needs demand. For instance, some vehicles might emphasize high semantic accuracy and tolerate some delays, while others may prioritize a higher semantic transmission rate even if it slightly compromises accuracy. To capture these varying preferences, we propose a QoE model that integrates both semantic accuracy and semantic rate. The QoE for vehicle $q$ in platoon $n$ is defined as:
$\text{QoE}_{q}^{n}$ $=\sum_{q\in Q}\left[\omega_{q}\text{Score}_{R}(\phi_{q})+(1-\omega_{q})\text{Score}_{A}(\xi_{q})\right]$ (19)
$=\sum_{q\in Q}\left[\frac{\omega_{q}}{1+e^{\gamma_{q}(\phi_{target}-\phi_{q})}}+\frac{1-\omega_{q}}{1+e^{\delta_{\mu}(\xi_{target}-\xi_{q})}}\right],$
where $\text{QoE}_{q}^{n}$ represents the quality of experience for vehicle $q$ in platoon $n$. $\omega_{q}$ is the weight assigned to the semantic rate for vehicle $q$, while $1-\omega_{q}$ represents the weight for semantic accuracy. $\phi_{target}$ and $\xi_{target}$ are the target levels of semantic rate and accuracy considered optimal for the platoon’s safe operation, and $\gamma_{q}$ and $\delta_{\mu}$ are parameters that determine the sensitivity of the scoring functions to deviations from these target values.
@@ II-C Optimization Problem
In this section, we formulate the semantic-aware resource allocation problem with the dual objectives of maximizing the quality of experience (QoE) for all vehicles within the platooning network and enhancing success rate of effective semantic information transmission (SRS) of V2V. In the optimization problem, we have the variables such as channel assignment, power allocation, and the selection of the number of transmitted semantic symbols for both text and image data. The problem can be formally defined as follows:
$\begin{array}[]{l}\mathop{{\rm{Maximize}}}\limits_{{\beta_{n}},{\rho_{n}},{p_{n}},{u_{m}}}\\ \sum\limits_{n\in N}{\sum\limits_{m\in{M_{n}}}{\left[{QoE_{m}^{n}{\rm{+}}\lambda\Pr\left({\sum\limits_{t=1}^{T}{\sum\limits_{m=1}^{M}{{\beta_{n,k}}}}\phi_{k}^{d}[m,t]\geq\frac{{{B_{s}}}}{{\Delta T}}}\right)}\right]}}\end{array}$ (20a)
@@
$\sum_{n\in N}\beta_{n,k}\leq 1,\quad\forall k\in W,$ (20c)
$\sum_{k\in W}\beta_{n,k}\leq 1,\quad\forall n\in N,$ (20d)
$u_{m}^{T}\in{0,1,\dots,u_{m,\max}^{T}},\quad\forall m\in M_{n},$ (20e)
$u_{m}^{I}\in{0,1,\dots,u_{m,\max}^{I}},\quad\forall m\in M_{n},$ (20f)
$0\leq p_{m}\leq p_{\max},\quad\forall m\in M_{n},\forall k\in W,$ (20g)
$Scor{e_{R,m}},Scor{e_{A,m}}\geq{G_{th}},\quad\forall m\in{M_{n}},\forall k\in W,$ (20h)
$0\leq\sum_{k=1}^{K}\sum_{n=1}^{N}\sum_{m=1}^{M}B_{s}\leq B_{s}^{\max}.$ (20i)
The objective function in (20a) aims to maximize the combined QoE of all vehicles and SRS within the time budget ${\Delta T}$. Here, $\lambda$ is a weight factor that balances the importance of QoE and timely delivery of safety-critical messages. The term $\Pr\left(\cdot\right)$ represents the probability the semantic packets are successfully transmitted over the V2V link in time. Constraints (20b) ensure that the channel allocation variable ${\beta_{n,k}}$ to be binary, meaning each subchannel $k$ is either assigned to platoon $n$ or not. (20c) imposes each subchannel to be allocated to at most one platoon, while (20d) ensures each platoon can be assigned at most one subchannel. Constraints (20e) and (20f) define the range for the number of semantic symbols transmitted for both text ${u_{m}^{T}}$ and image ${u_{m}^{I}}$ data. Constraint (20g) constrains the transmission power ${p_{m}}$ for each vehicle, within system boundaries a space. Constraints (22) ensure that the semantic rate and accuracy scores meet the minimum threshold ${G_{th}}$, guaranteeing acceptable QoE for all users. Finally, constraint (20i) ensures that the total semantic information size ${{B_{s}}}$ transmitted over V2V links is within the permissible limit $B_{s}^{\max}$, representing the upper bound of the semantic load capacity.
To address the problem, we model it as a MDP suitable for DRL. The optimization challenge in dynamic vehicular networks is complex and non-convex due to rapidly changing topologies, link interference, and varying communication demands. These aspects make traditional optimization methods ineffective. DRL is well-suited for this context as it learns adaptive policies through interaction with the environment. However, centralized DRL approaches suffer from high communication overhead and scalability issues. We therefore utilize MADDPG, a decentralized learning framework that enables each vehicle (agent) to independently learn optimal strategies while considering the actions of others. This is essential for multi-agent systems like platooning. Additionally, MADDPG’s capacity to manage continuous action spaces aligns it well with the complex resource management tasks in V2X communication.
@@ III Proposed SAMRAMARL Algorithm Approach
$\beta_{n,k}^{t}\in[0,1],\quad\forall n\in N,\forall k\in W.$ (21)
This relaxation transforms the problem from a mixed-integer non-linear programming (MINLP) problem into a non-linear programming (NLP) problem, which is more amenable to continuous optimization techniques. After solving the relaxed problem, a thresholding method can be used to map the continuous values back to binary decisions, providing a practical solution to the original problem. Additionally, the objective function (20a), which is a weighted sum of the QoE term and the probability of timely delivery, is further simplified to facilitate computation. For the success rate of semantic information transmission (SRS), we approximate it using a logistic function:
$\frac{1}{{1+\exp(-\alpha(\sum\limits_{t=1}^{T}{\sum\limits_{m=1}^{M}{{\beta_{n,k}}}}\phi_{k}^{d}[m,t]-\frac{{{B_{s}}}}{{\Delta T}}))}},$ (22)

### v2

@@ II-B Novel Metrics for Resource Management in platooning Networks
In vehicular platooning networks, effective resource management is crucial for maintaining efficient communication, particularly for autonomous driving applications. To optimize resource allocation and improve overall system performance, we propose novel metrics (semantic similarity and semantic rate) and quality of experience (QoE). These metrics aim to assess the efficiency of semantic communication by directly quantifying the amount of meaningful information being transmitted, going beyond traditional metrics that focus mainly on bit-level accuracy.
Semantic similarity is denoted as the extent to which the semantic information recovered at the receiver corresponds to the original information. Semantic rate is defined as the amount of semantic information transmitted per unit time. Both metrics are essential in vehicular networks, where communication must be both accurate and timely to ensure safe and coordinated actions within a platoon. Semantic rate is modeled based on semantic entropy. Unlike Shannon entropy, which measures the uncertainty of a source in terms of bits, semantic entropy quantifies the amount of meaningful information that a source carries with respect to a specific task. Calculating exact semantic entropy can be intractable for complex tasks. As such, we adopt deep learning (DL) models, such as DeepSC for single-modal and MU-DeepSC for multi-modal tasks, to approximate the semantic entropy[34]. This approximation, expressed in semantic units (suts), serves as the basis for optimizing resource allocation in platooning networks, where the goal is to transmit only the most critical feature information.
However, changes in the length of transmitted semantic symbols have opposite effects on semantic similarity and rate. As the length of transmitted semantic symbols decreases, the rate increases, but the similarity decreases. Therefore, we introduce a QoE model that assesses the overall performance of semantic communication systems by combining semantic accuracy and semantic rate. The QoE model proposed allows users to independently adjust their preferences for rate and accuracy, thus explicitly managing the trade-off between them.
II-B1 Semantic Similarity
For single-modal text communication, we employ semantic similarity ${\xi_{SM}}$) as a performance metric to evaluate the accuracy of semantic transmission [32]. This metric quantifies how closely the received message resembles the transmitted message in terms of its meaning. The semantic similarity can be expressed as:
@@ II-B1 Semantic Similarity
where the function ${\Psi_{SM}}$ maps these inputs to the semantic similarity score, and $0\leq{\xi_{SM}}\leq 1$, with ${\xi_{SM}}=1$ indicating the highest similarity between the two sentences and ${\xi_{SM}}=0$ implying no similarity. The term ${u_{m}^{T}}$ represents the average number of symbols required to encode each word as a semantic symbol (i.e., the semantic symbol length) by vehicle $m$, in units of suts/word. $SIN{R_{T,n,m/BS}}\left[k\right]$ denotes the SINR for text communication received at vehicle $m$ in platoon $n$ or at the BS over subchannel $k$.
For multi-modal communication, such as the fusion of text and image data, the semantic similarity $\xi_{MM}$ assesses joint performance of them. This metric evaluates how well the combined text and image information matches the expected outcome. The semantic similarity in this context is given by:
@@
where the function ${\Psi_{MM}}$ maps these inputs to the overall semantic similarity score. ${u_{m}^{T}}$ and ${u_{m}^{I}}$ denote the semantic symbol length for text and image, respectively, by vehicle $m$.
II-B2 Semantic Rate
The semantic rate measures the volume of meaningful feature information transmitted by a vehicle per second, reflecting the effectiveness of data communication in vehicular networks. Unlike traditional data rates, that focus on raw data transfer, the semantic rate emphasizes the transmission of information that is relevant for understanding and decision-making processes in specific tasks. The semantic rate ${\phi_{n}}$ is defined as:
${\phi_{m}}=\frac{{W{{\tilde{H}}_{SM}}}}{{u_{m}^{T}}},$ (15)
where $W$ represents the bandwidth allocated to the communication channel, ${{{\tilde{H}}_{SM}}}$ represents the approximate semantic entropy of the textual data. Semantic entropy is a critical concept that measures the amount of meaningful content within a message that is necessary for a specific task, rather than merely quantifying the raw data. Semantic entropy $H(X;Y)$ is defined as the minimum expected number of semantic symbols required to accurately predict or infer a task $Y$ from the data $X$ [35].
$\begin{array}[]{l}H(X;Y)\triangleq\mathop{\min}\limits_{{{\rm E}_{s}}}{\rm E}(\dim(Cod{e^{{{\rm E}_{s}}}}(X)))\\ s.t.\qquad{\rm{}}P(Y|Cod{e^{{{\rm E}_{s}}}}(X)){\rm{}}\approx P(Y|X),\end{array}$ (16)
where $Cod{e^{{{\rm E}_{s}}}}(X)$ represents the semantic symbols extracted from $X$ using a semantic encoder ${{{\rm E}_{s}}}$, and $P(Y|Cod{e^{{{\rm E}_{s}}}}(X))$ approximates the conditional probability of predicting $Y$ given $X$. This definition emphasizes that the goal of semantic entropy is to retain only the most relevant information for a particular task, optimizing communication efficiency, especially in dynamic and resource-constrained environments like vehicular networks. Computing an exact measure of semantic entropy is challenging due to the difficulty in determining an optimal semantic encoder. To address this, we use a deep learning-based approximation denoted as $\tilde{H}(X;Y)$.
$\begin{array}[]{l}\tilde{H}(X;Y)\triangleq\mathop{\min}\limits_{s{e_{\alpha}}}{\rm E}(\dim(Cod{e^{s{e_{\alpha}}}}(X)))\\ s.t.\qquad\left\lvert P(Y|Cod{e^{s{e_{\alpha}}}}(X)){\rm{}}-P(Y|X)\right\lvert\textless\varepsilon,\end{array}$ (17)
where the constraint indicates that the gap between $P(Y|Cod{e^{s{e_{\alpha}}}}(X))$ and $P(Y|X)$ can not exceed $\varepsilon$. This approach allows us to estimate the semantic content of a message by employing a well-trained semantic encoder ${s{e_{\alpha}}}$, capturing the key elements of the original data $X$ that are relevant for predicting the task $Y$.
@@ II-B2 Semantic Rate
${\phi_{{m_{{}_{T}}}}}=\frac{{W\tilde{H}_{MM}^{T}}}{{u_{m}^{T}}},{\rm{}}{\phi_{{m_{{}_{I}}}}}=\frac{{W\tilde{H}_{MM}^{I}}}{{u_{m}^{I}}},$ (18)
@@
$\sum_{t=1}^{T}\sum_{m=1}^{M_{n}}{{\beta_{n,k}}}\phi_{k}[m,t]\geq\frac{B_{s}}{\Delta T},\quad\forall m\in\mathcal{M}_{n},\forall k\in\mathcal{K},$ (19)
where $B_{s}$ denotes the size of the periodically generated V2V payload in semantic units (suts), $\Delta T$ is the channel coherence time, and $\phi_{k}[m,t]$ represents the semantic rate of the $k$-th V2V link for vehicle $m$ at time slot $t$. The variable ${{\beta_{n,k}}}$ indicates whether subchannel $k$ is assigned to platoon $n$ (i.e., ${{\beta_{n,k}}}=1$ if assigned, 00 otherwise). This condition ensures that the required amount of semantic information is delivered within the time constraints, thus enabling safe and synchronized movements within the platoon.
II-B3 QoE
Different vehicles within a platoon may prioritize these different metrics depending on their specific application needs demand. For instance, some vehicles might emphasize high semantic accuracy and tolerate some delays, while others may prioritize a higher semantic transmission rate even if it slightly compromises accuracy. To capture these varying preferences, we propose a QoE model that integrates both semantic accuracy and semantic rate. The QoE for vehicle $m$ in platoon $n$ is defined as:
$\text{QoE}_{m}^{n}$ $=\sum_{m\in\mathcal{M}_{n}}\left[\omega_{m}\text{Score}_{R}(\phi_{m})+(1-\omega_{m})\text{Score}_{A}(\xi_{m})\right]$ (20)
$=\sum_{m\in\mathcal{M}_{n}}\left[\frac{\omega_{m}}{1+e^{\gamma(\phi_{target}-\phi_{m})}}+\frac{1-\omega_{m}}{1+e^{\delta(\xi_{target}-\xi_{m})}}\right],$
where $\text{QoE}_{m}^{n}$ represents the quality of experience for vehicle $m$ in platoon $n$. $\omega_{m}$ is the weight assigned to the semantic rate for vehicle $m$, while $1-\omega_{m}$ represents the weight for semantic accuracy. $\phi_{target}$ and $\xi_{target}$ are the target levels of semantic rate and accuracy considered optimal for the platoon’s safe operation, and $\gamma$ and $\delta$ are parameters that determine the sensitivity of the scoring functions to deviations from these target values.
@@ II-C Optimization Problem
In this section, we formulate the semantic-aware resource allocation problem with the dual objectives of maximizing QoE for all vehicles within the platooning network and enhancing SRS of V2V. If only QoE is optimized, high QoE may be traded off by increasing the semantic symbol length to improve semantic similarity, but this will take up more channel resources and crowd out emergency message transmission opportunities for other vehicles, thus failing to transmit high-priority safety messages (e.g., emergency braking warnings) successfully within the strict time limit ($\Delta T$). In the autonomous vehicle platoon, the loss of latency-sensitive messages may lead to serious accidents. The proposed joint optimization framework bridges the above drawbacks.
In the optimization problem, we have the variables such as channel assignment, power allocation, and the selection of the length of transmitted semantic symbols for both text and image data. The problem can be formally defined as follows:
$\begin{array}[]{l}\mathop{{\rm{Maximize}}}\limits_{{\beta_{n}},{\rho_{n}},{p_{n}},{u_{m}^{T}},{u_{m}^{I}}}\\ \sum\limits_{n\in\mathcal{N}}{\sum\limits_{m\in\mathcal{M}_{n}}{\left[{QoE_{m}^{n}{\rm{+}}\lambda\Pr\left({\sum\limits_{t=1}^{T}{\sum\limits_{m=1}^{M_{n}}{{\beta_{n,k}}}}\phi_{k}[m,t]\geq\frac{{{B_{s}}}}{{\Delta T}}}\right)}\right]}}\end{array}$ (21a)
@@
$\sum_{n\in\mathcal{N}}\beta_{n,k}\leq 1,\quad\forall k\in\mathcal{K},$ (21c)
$\sum_{k\in\mathcal{K}}\beta_{n,k}\leq 1,\quad\forall n\in\mathcal{N},$ (21d)
$0\leq u_{m}^{T}\leq u_{m,\max}^{T},u_{m}^{T}\in\mathbb{N},\quad\forall m\in\mathcal{M}_{n},$ (21e)
$0\leq u_{m}^{I}\leq u_{m,\max}^{I},u_{m}^{I}\in\mathbb{N},\quad\forall m\in\mathcal{M}_{n},$ (21f)
$0\leq p_{m}\leq p_{\max},\quad\forall m\in\mathcal{M}_{n},\forall k\in\mathcal{K},$ (21g)
$Scor{e_{R,m}},Scor{e_{A,m}}\geq{G_{th}},\quad\forall m\in\mathcal{M}_{n},\forall k\in\mathcal{K},$ (21h)
$0\leq\sum_{k=1}^{K}\sum_{n=1}^{N}B_{s}\leq B_{s}^{\max}.$ (21i)
The objective function in (21a) aims to maximize the combined QoE of all vehicles and SRS within the time budget ${\Delta T}$. Here, $\lambda$ is a weight factor that balances the importance of QoE and timely delivery of safety-critical messages. The term $\Pr\left(\cdot\right)$ represents the probability the semantic packets are successfully transmitted over the V2V link in time, denoted as SRS. Constraints (21b) ensure that the channel allocation variable ${\beta_{n,k}}$ to be binary, meaning each subchannel $k$ is either assigned to platoon $n$ or not. (21c) imposes each subchannel to be allocated to at most one platoon, while (21d) ensures each platoon can be assigned at most one subchannel. Constraints (21e) and (21f) define the range for the number of semantic symbols transmitted for both text ${u_{m}^{T}}$ and image ${u_{m}^{I}}$ data. Constraint (21g) constrains the transmission power ${p_{m}}$ for each vehicle, within system boundaries a space. Constraints (21h) ensure that the semantic rate and accuracy scores meet the minimum threshold ${G_{th}}$, guaranteeing acceptable QoE for all users. Finally, constraint (21i) ensures that the total semantic information size ${{B_{s}}}$ transmitted over V2V links is within the permissible limit $B_{s}^{\max}$, representing the upper bound of the semantic load capacity.
In the proposed optimization problem, discrete variables such as binary channel allocation variables and continuous variables such as transmit power coexist, and nonlinearity is introduced by sigmoid functions. This makes it a mixed integer nonlinear programming (MINLP) problem. Furthermore, the dynamic network topology and time-varying channel conditions increase the complexity. These aspects make traditional optimization methods ineffective. To address the problem, we model it as a Markov decision process (MDP) suitable for DRL. DRL is well-suited for this context as it learns adaptive policies through interaction with the environment[36]. However, centralized DRL approaches suffer from high communication overhead and scalability issues. We therefore utilize multi-agent deep deterministic policy gradient (MADDPG), a decentralized learning framework that enables each vehicle (i.e., agent) to independently learn optimal strategies while considering the actions of others. This is essential for multi-agent systems like platooning. Additionally, MADDPG’s capacity to manage continuous action spaces aligns it well with the complex resource management tasks in V2X communication.
@@ III Proposed SAMRAMARL Algorithm Approach
$\beta_{n,k}^{t}\in[0,1],\quad\forall n\in\mathcal{N},\forall k\in\mathcal{K}.$ (22)
This relaxation transforms the problem from a MINLP problem into a non-linear programming (NLP) problem, which is more amenable to continuous optimization techniques. After solving the relaxed problem, a thresholding method can be used to map the continuous values back to binary decisions, providing a practical solution to the original problem. Additionally, the objective function (21a), which is a weighted sum of the QoE term and the probability of timely delivery, is further simplified to facilitate computation. For SRS, we approximate it using a logistic function:
$\frac{1}{{1+\exp(-\alpha(\sum\limits_{t=1}^{T}{\sum\limits_{m=1}^{M_{n}}{{\beta_{n,k}}}}\phi_{k}[m,t]-\frac{{{B_{s}}}}{{\Delta T}}))}},$ (23)